Calculate the volume of oxygen at 15°c and 745mmhg which could be obtained by heating 10g of potassium trioxochlorate.
1
Expert's answer
2019-11-21T06:26:03-0500

"PV=nRT,"

Where P is pressure (745 mmhg), V the volume to be calculated, n is the number of moles , R is a constant which is usually 8.314 j/kg/ mol and T is the temperature(288.15K)

2KClO3 ==> 2KCl + 3O2

moles of kClO3 =10/122.55 which is 0.0816moles

number of moles of O2= 3/2*0.0816 which is 0.1224moles

Substituting in the formula"V=nRT\/P;"

v=0.1224*8.314*288.15/745

V=0.3936m3